Gold has a molar mass of 197 g/mol. (a) How many moles of gold are in a 1.67 g sample of pure gold? (b) How many atoms are in the sample?
Answer #1

1 mole has mass=197 grams

==>1 gram have number of moles=1/197

then 1.67 gram will have number of moles=1.67/197=0.0085

b)as 1 mole=6.023*10^23

number of atoms in the sample=6.023*10^23*0.0085=5.1195*10^21